Using your data from the experiment, calculate the initial moles of HCl that you started with. ( for this problem you will use your initial concentration of HCl and your 50 mL volume to solve).

Trial 1 trail 2
Mass of KHP (KC8H5O4) (g) .29 g .21 g
Moles of KHP .000422 .00422
Moles of NaOH .01M .01M
Volume of NaOH .0247 .0222
Molarity of NaOH .0575 .0464
Average molarity of NaOH .0516 M
Answer #1

Number of mole s= molarity * volume

Average volume of NaOH = 0.0247 +0.022

= 0.02345 mlNaOH

mole of NaOH = molarity * volume in L

= 0.1* 0.02345/1000

= 2.345*10^-6 mole NaOH

The recation between HCl and NaOH is as follows:

HCl + NaOH= NaCl+ H2O

Therefore 2.345*10^-6 mole NaOH reacts with 2.345*10^-6 mole of HCl

Molarity = number of moles / volume in L

Here volume of HCl= 50 ml

0.050 L

Thus

Molarity = 2.345*10^-6 mole HCl/0.050 L

= 4.69*10^-5 M